OpenStudy (anonymous):

integral of (3t-2)/(t+1)

OpenStudy (turingtest):

easiest way is probably\[u=t+1\implies t=u-1\]\[du=dt\]

OpenStudy (turingtest):

long division would also work

OpenStudy (anonymous):

I tried it with long division and it was pretty easy that way.

OpenStudy (anonymous):

i like the gimmick of writing \[3t-2=3t+3-5\]

OpenStudy (anonymous):

where do I go with long division after I get 3 remainder -5?

OpenStudy (anonymous):

so you get \[\frac{3t-2}{t+1}=\frac{3t+3-5}{t+1}=1-\frac{5}{t-1}\]

OpenStudy (anonymous):

oops \[\frac{3t-2}{t+1}=\frac{3t+3-5}{t+1}=3-\frac{5}{t-1}\]

OpenStudy (anonymous):

then integrate each piece first one gives you \(3x\) second gives you \(-5\ln(t-1)\)

OpenStudy (anonymous):

ah got it thanks very much

OpenStudy (anonymous):

damn another typo!! should be \[3x-5\ln(x+1)\]

OpenStudy (anonymous):